OpenStudy (anonymous):

Determine any restrictions on the variable. x^2+2x/x^2+4

OpenStudy (solomonzelman):

set the denominator equal to ZERO. and solve for x.

OpenStudy (anonymous):

Restrictions are just when your denominator = 0. So yes, do what @SolomonZelman says.

OpenStudy (anonymous):

x^2=-4?

OpenStudy (anonymous):

it is always the case that \(x^2\geq 0\) and so \(x^2+4\geq 4\) i.e. it is never zero, so no restrictions